OpenStudy (anonymous):

I need a detailed answer please integrate 1/((9-4x^2)^(-1/2))

OpenStudy (across):

\[\int\frac{dx}{(9-4x^2)^{-\frac{1}{2}}}.\]Is this what you're trying to integrate?

OpenStudy (anonymous):

\[\int\limits 1\div(\sqrt(9-4x^2))\] this is what i am trying to integrate. Thank you

OpenStudy (across):

\[\int\frac{dx}{(9-4x^2)^{\frac{1}{2}}}.\]This then?

OpenStudy (anonymous):

yes perfect